Solve for triangle missing side b, missing bottom a, other side 6 with 2 angles both @ 45 degrees. Thank you
Answer by ReadingBoosters(3246)

The 2 angles at 45 degrees is the key hint for the right triangle with a 1,1 sqrt(2) pattern.
Knowing a^2+b^2=c^2, then c^2 = 36
a and b are the same, so
a^2 + a^2 = 36
2a^2 = 36
divide both sides by 2
a^2 = 18
a = sqrt(18)
b = sqrt(18)
c = 6
......................
Delighted to help
-Reading Boosters
We want for others what we want for ourselves.